OpenStudy (anonymous):

A few math 8th grade problems Please help me. PLEASE! Solve the system of equations. Use the substitution method to solve. 4x – 2y = 24 x – y = 3

OpenStudy (ivettef365):

Since you have x - y = 3 you can say that x = y+3 then substitute on first equation the x for (y+3), to find value of y

OpenStudy (anonymous):

Ok?

OpenStudy (anonymous):

What next then?

OpenStudy (anonymous):

Ive, what next?

OpenStudy (texaschic101):

x - y = 3 ---> x = 3 - y now sub 3 - y in for x in the other equation 4x - 2y = 24 4(3 - y) - 2y = 24 can you take it from here ?